OpenStudy (anonymous):

What is the missing exponent? (6^__)-4 = 6^20

OpenStudy (anonymous):

@arabpride @Awesome781 @blah124 @babbygirl12345 @hmundy @Firejay5 @zimmah @zylaai @Luigi0210 @LunyMoony @o0Ollie0o @protected @Paulywood16 @cwrw238 @CrazyTurtle483 @Sheraz12345 @sammixboo

OpenStudy (anonymous):

x * -4 = 20

OpenStudy (anonymous):

since \[\LARGE (n^p)^q=n^{p*q}\]

OpenStudy (anonymous):

5

OpenStudy (anonymous):

not quite

OpenStudy (anonymous):

-5 xD?

OpenStudy (anonymous):

that'd make -20

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

What is the value of the expression?

OpenStudy (anonymous):

If that is a new question it needs to be in a new thread :p

OpenStudy (anonymous):

\[\Large n^p * n^q = n^{p+q}\] \[\Large \frac{ n^p }{ n^q }=n^{p-q}\]

OpenStudy (anonymous):

so \(\Large 7^{-3+9-(-8)}\)